The perimeter of a rectangle is 86cm. Twice the width exceeds the length by 2cms. Find the dimensions. I used this equation P=2l+2W, but I don't know how to combine the terms
Answer by LilSkittleMd(119)

P=2L+2W
The second equation is 2W=L+2 (Twice the width exceeds the length by 2)
You need to put this equation into a "L equals" or "W equals" so you can plug it into the P equation. Use an L equation, it is easier to use. Just subtract 2 from both sides.
2W-2=L
Now just put everything together
86=2(2W-2)+2W
86=4W-4+2W
86=6W-4
90=6W
15=W
L=2(15)-2
L=30-2
L=28
The width is 15 cm and the length is 28 cm